# Robustness-Congruent Adversarial Training for Secure Machine Learning Model Updates

### v1

@@ 3.1.1 Solution Algorithm
Under these premises, we can finally state the RCAT algorithm used to optimize Eq. (8), given as Algorithm 1. The algorithm starts by initializing the model $f_{\boldsymbol{w}}$ with $f^{\rm src}$ (Algorithm 1). It then runs for $E$ epochs (Algorithm 1), looping over the whole training samples in each epoch (Algorithm 1). In each iteration, RCAT runs the attack algorithm $a$ to optimize the adversarial example $\boldsymbol{x}^{\star}$ within the feasible domain $\mathcal{B}(\boldsymbol{x})$ (Algorithm 1). Then, it uses the adversarial example $\boldsymbol{x}^{\star}$ to update the model parameters $\boldsymbol{w}$ along the gradient direction (Algorithm 1). While we present here a sample-wise version of the RCAT algorithm, it is worth remarking that it is straightforward to implement it using in batch-wise manner.
@@ 3.2 Consistency Results
We now analyze the formal guarantees that can be derived for the problem of learning when including a non-regression constraint aimed to reduce regression of accuracy and robustness. Under the assumption that samples are independent and identically distributed (i.i.d.) and in the absence of constraints, it is well known that learning algorithms that optimize either the accuracy or robustness yield consistent statistical estimators, i.e., they converge to the correct model as the number of samples $n$ increases, with a rate of $O(\nicefrac{{1}}{{\sqrt{n}}})$ in the general case [9, 10]. This means that collecting samples to increase the training set is generally worth, as it is expected to increase the performance of the model. However, neither consistency nor the convergence rate are guaranteed when the optimization problem includes constraints [9]. In this section, we are the first to show that the inclusion of the non-regression constraint inside the optimization problem produces an estimator that is consistent on both accuracy and robustness, while also preserving the same convergence rate.
@@ 3.2.1 Preliminaries
Let us conclude this section by discussing the role of $\hat{\epsilon}$ and $\epsilon$ in the aforementioned constraints. While the desired $\epsilon\in[0,\infty)$ should be set to zero to have zero regression (i.e., no negative flips of accuracy or robustness), $\hat{\epsilon}\in[0,\infty)$ is usually set to a small value $\hat{\epsilon}>0$ for two main reasons. Theoretically, as discussed in the following, this is a sufficient condition to ensure Problem (14) to be consistent with respect to Problem (15). Practically, as shown in Sect. 4, setting $\hat{\epsilon}>0$ allows us to obtain larger improvements in the test error of $\hat{f}$ since $\mathcal{F}$ may be not perfectly designed and moreover the number of (noisy) samples is limited.
@@ 3.2.3 Consistency of PCT, PCAT, and RCAT
The previous section proves consistency of $\hat{f}$ with the usual convergence rate when the learning problem includes a non-regression constraint. This means that designing updates that reduce NFs or RNFs does not compromise any of the standard properties that normally hold for learning algorithms.
@@
It is worth remarking that not only PCAT and RCAT provide consistent statistical estimators, but that this also applies to PCT [1] and to any algorithm derived from a formulation that includes a non-regression constraint or penalty term.
@@ 4 Experimental Analysis
To show the validity of the RCAT method, we report here an extensive experimental analysis involving several robust machine-learning models designed for image classification. After describing the experimental setup (Sect. 4.1), we demonstrate that the problem of regression of robustness is relevant when replacing a robust machine-learning model with an improved state-of-the-art model (Sect. 4.2). We then show that RCAT can better mitigate the regression compared to PCT, PCAT, and naïve model update strategies (Sect. 4.3).
@@ 4.1 Experimental Setup
We detail here the experimental setup used in our analyses.
@@
Dataset. We consider the CIFAR-10 dataset, as it has been extensively used to benchmark the adversarial robustness of image classifiers in the well-known RobustBench framework [2]. CIFAR-10 consists of 60,000 color images of size $32\times 32$ that belong to $10$ different classes, including 50,000 training images and 10,000 test images. In our experiments, we use 40,000 training images to update models with PCT, PCAT, and RCAT, and the remaining 10,000 training images as a validation set to tune the hyperparameters of the aforementioned methods. The performance metrics (detailed below) are then evaluated on a subset of 2,000 test images.
@@
Robust Models. We consider seven CIFAR-10 robust models from RobustBench [2], denoted from the least to the most robust with $M_{1},\ldots,M_{7}$, and originally proposed respectively in [21, 22, 23, 24, 25, 26, 27]. These models are evaluated in RobustBench [2] against $\ell_{\infty}$-norm attacks with a perturbation budget of $\epsilon=8/255$. They use different architectures, training strategies, and exhibit different trade-offs between accuracy and robustness, enabling us to simulate different scenarios in which there would be a clear incentive in replacing an old model with an improved one.
@@
Model Updates. We consider 4 different model update strategies: (i) naïve, where we just replace the old model with an already-trained, more robust one; (ii) PCT; (iii) PCAT; and (iv) RCAT. When using PCT, PCAT, and RCAT, we fine-tune models using E = 12 training epochs, and batches of 500 samples. We set the learning rate $\eta=$ $10^{-3}$. PCAT and RCAT also require implementing adversarial training using an attack algorithm $a$, as detailed in Algorithm 1. To this end, we follow the implementation of the Fast Adversarial Training (FAT) approach proposed in [28], which has been shown to provide similar results to the computationally-demanding adversarial training (AT) proposed in [3], while being far more efficient. The underlying idea of FAT is to randomly perturb the initial training samples and then use a fast, non-iterative attack to compute the corresponding adversarial perturbations. In particular, instead of using the iterative Projected Gradient Descent (PGD) attack as done in AT [3], FAT uses the so-called Fast Gradient Sign Method (FGSM) [5], i.e., a much faster, non-iterative $\ell_{\infty}$-norm attack. Even if FGSM is typically less effective than PGD in finding adversarial examples, when combined with random initialization to implement adversarial training, it turns out to achieve competing results [28].
@@
Hyperparameter Tuning. We choose the hyperparameters of PCT, PCAT, and RCAT that minimize the overall number of negative flips (NFs) and robustness negative flips (RNFs) on the validation set, to achieve a reasonable trade-off between reducing the regression of accuracy and that of robustness with respect to the naïve strategy of replacing the previous model with the more recent one. For PCT (Eq. 2) and PCAT (Eq. 7) baseline methods, we fix $\lambda=1$ and $\alpha=1$, and we run a grid-search on $\beta\in\{1,2,5,10\}$, as recommended by Yan et al. [1]. For RCAT, we run a grid-search on $(\alpha,\beta)\in\{(0.75,0.2),(0.7,0.2),(0.5,0.4),(0.3,0.6)\}$ while fixing $\gamma=1-\alpha-\beta$. These configurations attempt to give more or less importance to the non-regression penalty term, while RCAT also re-balances the other loss components.
@@
Figure 2: Regression of robustness and accuracy for $M_{1},\ldots,M_{7}$ [21, 22, 23, 24, 25, 26, 27]. (a) Left: Robust error (%) of each model, sorted in descending order. Right: RNFs (%) attained when replacing old models (in rows) with new ones (in columns). (b) Left: Test error (%) of each model, sorted in descending order. Right: NFs (%) attained when replacing old models (in rows) with new ones (in columns). Values in the upper (lower) triangular matrices evaluate regression when the new model has better (worse) average robustness/accuracy than the old model.
Performance Metrics. We consider four relevant metrics, evaluated on $2,000$ test samples, to assess the performance of the given methods: (i) the test error, i.e., the percentage of misclassified (clean) test samples; (ii) the robust error, i.e., the percentage of misclassified adversarial examples; (iii) the fraction of negative flips (NFs, Eq. 1), to quantify the regression of accuracy; and (iv) the fraction of robustness negative flips (RNFs, Eq. 6), to evaluate the regression of robustness. To evaluate the robust error and the RNF rate, we optimize adversarial examples against each model using the AutoPGD attack [8] implemented in adversarial-library, [^4] Available at https://github.com/jeromerony/adversarial-library. with the Difference-of-Logit-Ratio (DLR) loss and $N=50$ iterations.
@@ 4.2 Evaluating Regression of Robustness
The first experiment presented here aims to empirically show that, once updated, machine-learning models are affected not only by a regression of accuracy, as shown in [1], but also by a regression of robustness. To show this novel phenomenon, as described in Sect. 4.1, we first rank the models $M_{1},...,M_{7}$ from the least to the most robust, based on the robust error evaluated on the $2,000$ test samples perturbed with AutoPGD. The models, along with their robust errors, are reported in order in 1(a) (left). Then, assuming a naïve update strategy that just replaces the old model with the new one, without performing any fine-tuning of the latter, we evaluate regression of robustness for all possible model pairs, as shown in 1(a) (right). Each cell of this matrix reports the regression of robustness (i.e., the fraction of RNFs) induced by the new model (reported in the corresponding column) when replacing the old one (reported in the corresponding row). The upper (lower) triangular matrix represents the cases in which the new model has better (lower) average robust error than the old one. The diagonal corresponds to replacing the model with itself, and thus no regression is observed. Even in the ideal case in which machine-learning models are updated with more robust ones (i.e., in the upper triangular matrix), we observe an RNF rate ranging from roughly $2\%$ to $6\%$. This implies that an increment in average robustness does not guarantee that previously-ineffective adversarial examples are still correctly classified.
We perform the same analysis on $M_{1},\ldots,M_{7}$ to quantify also their regression of accuracy. We thus re-order the models according to their reported test errors, as shown in 1(b) (left). We then consider again all the possible model pairs to simulate different model updates, and report the NFs corresponding to each naïve model update (i.e., by replacement) in 1(b) (right). Not surprisingly, all updates induce a regression of accuracy, even when the new models are more accurate than the old ones (i.e., in the upper triangular matrix), as also already shown by Yan et al. [1].
@@
TABLE I: Results for PCT, PCAT, and RCAT on the $10$ model updates with the highest RNF rates. In each update, denoted with ($M_{i},M_{j}$), the baseline model $M_{i}$ is replaced with $M_{j}$ using the naïve (replacement) strategy, or by fine-tuning $M_{j}$ with PCT, PCAT, and RCAT. For each update, we evaluate the test error, NFs, robust error, and RNFs (in %), and highlight in bold the method that achieves the best tradeoff between NFs and RNFs, i.e., the lowest value of their sum.
@@ 4.3 Reducing Regression in Robust Image Classifiers
After having quantified the non-negligible impact of RNFs in model updates, we show here how RCAT can tackle this issue, outperforming the competing model update strategies of PCT and PCAT. To this end, we first select suitable model pairs that enable simulating updates in which the new model improves both accuracy and robustness w.r.t. the old one. This amounts to considering only $14$ model pairs (out of the overall $42$). Among these $14$ cases, we exclude $4$ model pairs exhibiting an RNF rate lower than $3\%$, which results in retaining the $10$ cases with the highest RNFs listed in Table I. Let us also recall that, when using the naïve update strategy, we just replace the old model $M_{i}$ with the new model $M_{j}$, and measure the corresponding NFs and RNFs. When using PCT, PCAT, and RCAT, we initialize the new model to be fine-tuned with $M_{j}$. Finally, for RCAT, we set $f^{\rm src}=M_{j}$.
Results for Model Updates. From the results in Table I, it is clear that RCAT provides a better trade-off between the reduction of NFs and that of RNFs w.r.t. the competing approaches in almost all cases. RCAT indeed achieves lower values of the sum of NFs and RNFs, while PCT and PCAT mostly reduce NFs by improving accuracy at the expense of compromising robustness. More specifically, PCT almost always entirely compromises robustness, as it is not designed to preserve it after update. A paradigmatic example can be found in row 5 of Table I, where PCT recovers almost 10% of test error w.r.t. the naïve strategy, lowering the NFs down to  1%, but increasing the robust error by  43%, with obvious consequences for the RNF rate that increases by  35%.
Measuring Performance Improvements ($\Delta$-metrics). To better highlight the differences among PCT, PCAT, and RCAT, for each update ($M_{i}$, $M_{j}$), we compute their performance improvements w.r.t. the baseline naïve strategy using these four additional metrics, referred to as $\Delta$-metrics in the following:
@@
Analysis with $\Delta$-metrics. The results are reported in Fig. 3 and in Table II, along with the mean values of the $\Delta$-metrics, averaged on the model updates listed in Table I. From this analysis, the trends followed by PCT, PCAT, and RCAT become clearer: (i) PCT provides the largest improvement of accuracy (+4.62% on average) and reduction of NFs (+1.58% on average), at the expense of compromising adversarial robustness (-15.14% on average) and RNFs (-11.05% on average); (ii) PCAT works slightly better, but it still tends to improve accuracy (+2.59% on average) without retaining robustness (-4.38% on average), thereby improving NFs (+1.12% on average) and worsening RNFs (-2.16% on average); and (iii) RCAT finds a better accuracy-robustness trade-off than the competing approaches. In particular, it only slightly improves accuracy (+1.62% on average) and NFs (+1% on average), without affecting robustness (only -0.29% on average) and improving RNFs (+0.18% on average).
To summarize, performing model updates that retain high accuracy and robustness with low regression of both metrics is challenging. PCT and PCAT tend to recover accuracy but lose robustness. RCAT finds a better accuracy-robustness trade-off, providing a smaller improvement in accuracy and preserving robustness and, at the same time, minimizing the sum of NFs and RNFs, thereby effectively reducing both regression of accuracy and robustness.
Figure 3: Performance improvements ($\Delta$-metrics) for each model update, when comparing naïve model replacement with PCT, PCAT, and RCAT. Positive values represent improvements w.r.t. naïve updates. Left: this plot shows $\Delta$ Test Error vs $\Delta$ Robust Error, highlighting the trade-off between accuracy and robustness of each method. Right: this plot shows $\Delta$ NFs vs $\Delta$ RNFs, highlighting the trade-off between reducing NFs and RNFs of each method. For each method, we report the individual model update values (transparent points) from Table I, and the corresponding average (opaque point) from Table II.
TABLE II: Mean values of the $\Delta$-metrics, averaged on all model updates shown in Table I. Positive values mean better results, i.e., lower test/robust error and NF/RNF rate w.r.t. to replacing the old model with the new one (naïve updates).
$\Delta$ Test Error $\Delta$ NFs $\Delta$ Robust Error $\Delta$ RNFs
PCT +4.62 +1.58 -15.14 -11.05

### v2

@@ 3.1.1 Solution Algorithm
Under these premises, we can finally state the RCAT algorithm used to optimize Eq. (8), given as Algorithm 1. The algorithm starts by initializing the model $f_{\boldsymbol{w}}$ with $f^{\rm src}$ (Algorithm 1). It then runs for $E$ epochs (Algorithm 1), looping over the whole training samples in each epoch (Algorithm 1). In each iteration, RCAT runs the attack algorithm $a$ to optimize the adversarial example $\boldsymbol{x}^{\star}$ within the feasible domain $\mathcal{B}(\boldsymbol{x})$ (Algorithm 1). Then, it uses the adversarial example $\boldsymbol{x}^{\star}$ to update the model parameters $\boldsymbol{w}$ along the gradient direction (Algorithm 1). While we present here a sample-wise version of the RCAT algorithm, it is worth remarking that it is straightforward to implement it batch-wise.
@@ 3.2 Consistency Results
We now analyze the formal guarantees that can be derived for the learning problem when including a non-regression constraint to reduce regression of accuracy and robustness. Under the assumption that samples are independent and identically distributed (i.i.d.) and in the absence of constraints, it is well known that learning algorithms that optimize either the accuracy or robustness yield consistent statistical estimators, i.e., they converge to the correct model as the number of samples $n$ increases, with a rate of $O(\nicefrac{{1}}{{\sqrt{n}}})$ in the general case [9, 10]. This means that collecting samples to increase the training set is generally worthwhile, as it is expected to improve the performance of the model. However, neither consistency nor the convergence rate is guaranteed when the optimization problem includes constraints [9]. In this section, we are the first to show that including the non-regression constraint inside the optimization problem produces an estimator that is consistent on both accuracy and robustness, while also preserving the same convergence rate.
@@ 3.2.1 Preliminaries
Let us conclude this section by discussing the role of $\hat{\epsilon}$ and $\epsilon$ in the aforementioned constraints. While the desired $\epsilon\in[0,\infty)$ should be set to zero to have zero regression (i.e., no negative flips of accuracy or robustness), $\hat{\epsilon}\in[0,\infty)$ is usually set to a small value $\hat{\epsilon}>0$ for two main reasons. Theoretically, as discussed in the following, this is a sufficient condition to ensure Problem (14) to be consistent with respect to Problem (15). Practically, as shown in Sect. 4, setting $\hat{\epsilon}>0$ allows us to obtain larger improvements in the test error of $\hat{f}$ since $\mathcal{F}$ may be not perfectly designed and the number of (noisy) samples is limited.
@@ 3.2.3 Consistency of PCT, PCAT, and RCAT
The previous section proves the consistency of $\hat{f}$ with the usual convergence rate when the learning problem includes a non-regression constraint. This means that designing updates that reduce NFs or RNFs does not compromise any of the standard properties that normally hold for learning algorithms.
@@
It is worth remarking that not only PCAT and RCAT provide consistent statistical estimators, but this also applies to PCT [1] and any algorithm derived from a formulation that includes a non-regression constraint or penalty term.
@@ 4 Experimental Analysis
To show the validity of the RCAT method, we report here an extensive experimental analysis involving several robust machine-learning models designed for image classification. After describing the experimental setup (Sect. 4.1), we demonstrate that the problem of regression of robustness is relevant when replacing a robust machine-learning model with an improved state-of-the-art model (Sect. 4.2). We then show that RCAT can better mitigate the regression compared to PCT, PCAT, and naïve model update strategies, not only when considering a single update (Sect. 4.3), but also in the case of multiple sequential updates (Sect. 4.4).
@@ 4.1 Experimental Setup
Here, we detail the experimental setup used in our analyses.
@@
Datasets. We consider CIFAR-10 and ImageNet as they have been extensively used to benchmark the adversarial robustness of image classifiers in the well-known RobustBench framework [4]. CIFAR-10 consists of 60,000 color images of size $32\times 32$ belonging to $10$ different classes, including 50,000 training images and 10,000 test images. Following current evaluation standards [4], we use 40,000 training images to update models with PCT, PCAT, and RCAT, the remaining 10,000 training images as a validation set for hyperparameter tuning, and a subset of 2,000 test images to evaluate the performance metrics (as detailed below). For ImageNet, we consider only its validation set (i.e., 50,000 color images of size $224\times 224$ divided into 1,000 classes), using 36,000 images to update the models, 9,000 as a validation set for hyperparameter tuning, and the remaining 5,000 for testing.
@@
Robust Models. In our experiments, we consider robust models from RobustBench [4], evaluated against $\ell_{\infty}$-norm attacks with a perturbation budget of $\epsilon=8/255$ and $\epsilon=4/255$ for CIFAR-10 and ImageNet, respectively. In particular, we consider seven CIFAR-10 robust models denoted from the least to the most robust with $C_{1},\ldots,C_{7}$, and originally proposed respectively in [21, 22, 23, 24, 25, 26, 27]. We then consider five ImageNet robust models, denoted from the least to the most robust with $I_{1},\ldots,I_{5}$, and originally proposed in [28, 21, 29] ($I_{1},I_{2},I_{3}$), and [30] ($I_{4},I_{5}$). All the considered models use different architectures, training strategies, and exhibit different trade-offs between accuracy and robustness, enabling us to simulate different scenarios in which there would be a clear incentive in replacing an old model with an improved one.
@@
Model Updates. We consider four different model update strategies: (i) naïve, where we just replace the old model with an already-trained, more robust one; (ii) PCT; (iii) PCAT; and (iv) RCAT. When using PCT, PCAT, and RCAT, we fine-tune CIFAR-10 models using E = 12 training epochs, and batches of 500 samples. For ImageNet models, we use 10 epochs and batches of 128 samples for the first four models. We use batches of 64 samples for the last model instead, to fit in memory, and reduce the number of epochs to 5 to keep the number of iterations unchanged. We set the learning rate $\eta=$$10^{-3}$ for training on both datasets. PCAT and RCAT also require implementing adversarial training using an attack algorithm $a$, as detailed in Algorithm 1. To this end, we follow the implementation of the Fast Adversarial Training (FAT) approach proposed in [31], which has been shown to provide similar results to the computationally-demanding adversarial training (AT) proposed in [5], while being far more efficient. The underlying idea of FAT is to randomly perturb the initial training samples and then use a fast, non-iterative attack to compute the corresponding adversarial perturbations. In particular, instead of using the iterative Projected Gradient Descent (PGD) attack as done in AT [5], FAT uses the so-called Fast Gradient Sign Method (FGSM) [32], i.e., a much faster, non-iterative $\ell_{\infty}$-norm attack. Even if FGSM is typically less effective than PGD in finding adversarial examples, when combined with random initialization to implement adversarial training, it turns out to achieve competing results [31].
@@
Hyperparameter Tuning. We choose the hyperparameters of PCT, PCAT, and RCAT that minimize the overall number of negative flips (NFs) and robustness negative flips (RNFs) (i.e., their sum) on the validation set, to achieve a reasonable trade-off between reducing the regression of accuracy and that of robustness with respect to the naïve strategy of replacing the previous model with the more recent one. For PCT (Eq. 2) and PCAT (Eq. 7) baseline methods, we fix $\lambda=1$ and $\alpha=1$, and we run a grid-search on $\beta\in\{1,2,5,10\}$, as recommended by Yan et al. [1]. For RCAT, we run a grid-search on $(\alpha,\beta)\in\{(0.75,0.2),(0.7,0.2),(0.5,0.4),(0.3,0.6)\}$ while fixing $\gamma=1-\alpha-\beta$. These configurations attempt to give more or less importance to the non-regression penalty term, while RCAT also re-balances the other loss components.
@@
Figure 2: Regression of robustness and accuracy for the CIFAR-10 robust models $C_{1},\ldots,C_{7}$ [21, 22, 23, 24, 25, 26, 27]. (a) Left: Robust error (%) of each model, sorted in descending order. Right: RNFs (%) attained when replacing old models (in rows) with new ones (in columns). (b) Left: Test error (%) of each model, sorted in descending order. Right: NFs (%) attained when replacing old models (in rows) with new ones (in columns). Values in the upper (lower) triangular matrices evaluate regression when the new model has better (worse) average robustness/accuracy than the old model.
Performance Metrics. We consider four relevant metrics to evaluate the given methods: (i) the test error, i.e., the percentage of misclassified (clean) test samples; (ii) the robust error, i.e., the percentage of misclassified adversarial examples; (iii) the fraction of negative flips (NFs, Eq. 1), to quantify the regression of accuracy; and (iv) the fraction of robustness negative flips (RNFs, Eq. 6), to evaluate the regression of robustness. To evaluate the robust error and RNF rates, we optimize adversarial examples against each model using the AutoPGD [8] implementation from adversarial-library, [^4] Available at https://github.com/jeromerony/adversarial-library. with the Difference-of-Logit-Ratio (DLR) loss and $N=50$ iterations.
@@ 4.2 Evaluating Regression of Robustness
The first experiment presented here aims to empirically show that, once updated, machine-learning models are affected not only by a regression of accuracy, as shown in [1], but also by a regression of robustness. To show this novel phenomenon, as described in Sect. 4.1, we first rank the CIFAR-10 models $C_{1},...,C_{7}$ from the least to the most robust, based on the robust error evaluated on the $2,000$ test samples perturbed with AutoPGD. The models, along with their robust errors, are reported in order in 2(a) (left). Then, assuming a naïve update strategy that just replaces the old model with the new one, without performing any fine-tuning of the latter, we evaluate regression of robustness for all possible model pairs, as shown in 2(a) (right). Each cell of this matrix reports the regression of robustness (i.e., the fraction of RNFs) induced by the new model (reported in the corresponding column) when replacing the old one (reported in the corresponding row). The upper (lower) triangular matrix represents the cases where the new model has a better (lower) average robust error than the old one. The diagonal corresponds to replacing the model with itself; thus, no regression is observed. Even in the ideal case where machine-learning models are updated with more robust ones (i.e., in the upper triangular matrix), the RNF rate ranges from roughly $2\%$ to $6\%$. This implies that an increment in average robustness does not guarantee that previously-ineffective adversarial attacks on some samples remain unsuccessful after update.
We perform the same analysis on $C_{1},\ldots,C_{7}$ to quantify also their regression of accuracy. We thus re-order the models according to their reported test errors, as shown in 2(b) (left). We then consider again all the possible model pairs to simulate different model updates, and report the NFs corresponding to each naïve model update (i.e., by replacement) in 2(b) (right). Not surprisingly, all updates induce a regression of accuracy, even when the new models are more accurate than the old ones (i.e., in the upper triangular matrix), as also already shown by Yan et al. [1].
@@
TABLE I: Results for PCT, PCAT, and RCAT on the $10$ CIFAR-10 model updates with the highest RNF rates. In each update, denoted with ($C_{i},C_{j}$), the baseline model $C_{i}$ is replaced with $C_{j}$ using the naïve (replacement) strategy, or by fine-tuning $C_{j}$ with PCT, PCAT, and RCAT. For each update, we evaluate the test error, NFs, robust error, and RNFs (in %), and highlight in bold the method achieving the best trade-off between NFs and RNFs, i.e., the lowest value of their sum.
@@
TABLE II: Results for PCT, PCAT, and RCAT on the $4$ ImageNet model updates. In each update, denoted with ($I_{i},I_{j}$), the baseline model $I_{i}$ is replaced with $I_{j}$ using the naïve (replacement) strategy, or by fine-tuning $I_{j}$ with PCT, PCAT, and RCAT. For each update, we evaluate the test error, NFs, robust error, and RNFs (in %), and highlight in bold the method achieving the best trade-off between NFs and RNFs, i.e., the lowest value of their sum.
Test Error NFs Robust Error RNFs
($I_{1},I_{2}$) baseline 47.78 - 73.48 -
naïve 37.66 2.22 66.98 2.88
PCT 34.92 2.32 76.54 7.62
PCAT 38.26 2.66 65.38 2.32
RCAT 37.70 1.96 64.78 2.00
($I_{2},I_{3}$) baseline 37.66 - 66.98 -
naïve 31.22 2.84 57.22 2.14
PCT 28.42 2.94 83.52 17.82
PCAT 32.58 4.14 60.74 3.96
RCAT 30.90 2.24 55.88 1.62
($I_{3},I_{4}$) baseline 31.22 - 57.22 -
naïve 23.60 2.76 44.08 2.20
PCT 19.96 2.16 75.12 20.50
PCAT 24.68 2.28 44.02 1.62
RCAT 23.12 2.26 43.08 1.82
($I_{4},I_{5}$) baseline 23.60 - 44.08 -
naïve 22.12 2.74 40.64 2.54
PCT 17.70 2.00 80.58 37.22
PCAT 20.58 2.44 38.48 1.92
RCAT 21.24 2.30 39.54 2.06
@@ 4.3 Reducing Regression in Robust Image Classifiers
After quantifying the non-negligible impact of RNFs in model updates, we show how RCAT can tackle this issue, outperforming the competing model update strategies of PCT and PCAT. To this end, we first select suitable model pairs that enable simulating updates in which the new model improves both accuracy and robustness w.r.t. the old one. This amounts to considering only $14$ model pairs (out of the overall $42$) for CIFAR-10. Among these $14$ cases, we exclude $4$ model pairs exhibiting an RNF rate lower than $3\%$, which results in retaining the $10$ cases with the highest RNFs listed in Table I. Let us also recall that, when using the naïve update strategy, we just replace the old model $C_{i}$ with the new model $C_{j}$, and measure the corresponding NFs and RNFs. When using PCT, PCAT, and RCAT, we initialize the new model to be fine-tuned with $C_{j}$. For RCAT, we also set $f^{\rm src}=C_{j}$. For ImageNet models, we consider the four pairs reported in Table II, in which the new models always exhibit improved accuracy and robustness w.r.t. the old ones, with non-negligible NF and RNF rates of about 2-3%.
Results for Model Updates. The results in Table I and Table II report the performance metrics for each model update. [^5] We also estimate their standard deviation using the bootstrap method with 1,000 resamplings, and find that it is negligible, being approximately 0.6% for the test error, 0.25% for NFs, 1% for the robust error, and 0.5% for RNFs, for all methods, with minor variations. It is clear that RCAT provides a better trade-off between the reduction of NFs and that of RNFs w.r.t. the competing approaches in almost all cases. RCAT indeed achieves lower values of the sum of NFs and RNFs, while PCT and PCAT mostly reduce NFs by improving accuracy at the expense of compromising robustness. More specifically, PCT almost always entirely compromises robustness, as it is not designed to preserve it after update. A paradigmatic example can be found in row 5 of Table I, where PCT recovers almost 10% of test error w.r.t. the naïve strategy, lowering the NFs close to $1\%$, but increasing the robust error by approximately $43\%$ and the RNF rate by almost $35\%$. Similar trends are also reported for ImageNet models; e.g., considering the model update $(I_{4},I_{5})$, PCT reduces the test error by almost 5% while increasing the robust error and RNFs by approximately $40\%$ and $35\%$, respectively. Let us finally remark that we also compute the number of samples for which no adversarial example was found before update, but become misclassified after update, contributing to both NFs and RNFs as discussed in Sect. 2.2. These joint negative flips are always less than approximately 0.03% for CIFAR-10 and 0.2% for ImageNet, thus being negligible overall.
Measuring Performance Improvements ($\Delta$-metrics). To better highlight the differences among PCT, PCAT, and RCAT, for each update ($M_{i}$, $M_{j}$) (being $M_{i}$ and $M_{j}$ two generic CIFAR-10 or ImageNet models), we compute their performance improvements w.r.t. the baseline naïve strategy using the following four $\Delta$-metrics:
@@
Analysis with $\Delta$-metrics. The results are reported in Table III, along with the values of the $\Delta$-metrics averaged on the model updates listed in Table I and Table II. PCT provides the largest improvement in accuracy for both CIFAR-10 and ImageNet models, and a good reduction of NFs, at the expense of significantly worsening adversarial robustness (-15.14%/-22.50% for CIFAR-10 and ImageNet, respectively) and RNFs (-11.05%/-14.85%). PCAT works slightly better for CIFAR-10 models, improving accuracy and NFs by +2.59% and +1.12% on average, but decreasing the robust error and RNF rate by -4.38% and -2.16% on average. For ImageNet models, PCAT performs similarly to the naïve update strategy. The proposed RCAT method, instead, finds a better accuracy-robustness trade-off; in particular, it only slightly improves accuracy (+1.62%/+0.39% on average) and NFs (+1%/+0.55% on average) without significantly affecting robustness, while improving RNFs (+0.18%/+0.66% on average). We also note that for ImageNet models, RCAT is the best performing in terms of NFs (+0.55% on average). [^6] It is worth remarking here that, in general, the additional constraint of reducing NFs/RNFs in the considered update strategies may cause the test/robust error to increase (e.g., RCAT slightly worsens the robust error, on average, while reducing the RNF rate). This may become more evident in the case of more accurate/robust models.
TABLE III: Mean values of the $\Delta$-metrics, averaged on all CIFAR-10 (ImageNet) model updates in Table I (Table II). Positive values mean better results, i.e., lower test/robust error and NF/RNF rate w.r.t. to the naïve strategy (i.e., model replacement). For each $\Delta$-metric and dataset, we highlight in bold the highest improvement among the three methods.
$\Delta$ Test Err. $\Delta$ NFs $\Delta$ Rob. Err. $\Delta$ RNFs
CIFAR-10 PCT +4.62 +1.58 -15.14 -11.05
@@
ImageNet PCT +3.08 +0,30 -22.50 -14.85
PCAT +0.01 -0.17 +0.62 -0.03
RCAT +0.39 +0.55 +1.61 +0.66
4.4 Reducing Regression over Sequential Updates
We have considered so far a single-update setting in which the old model is replaced by the new one, using a given update method (being it naïve, PCT, PCAT, or RCAT). In this section, instead, we consider the case in which we perform multiple sequential updates to evaluate the behavior of the update strategies when used iteratively. To this end, in Table IV we report the results of RCAT, PCT, and PCAT when used to update the ImageNet models $I_{1},\ldots I_{5}$ sequentially. In this experiment, we set $(\alpha,\beta)=(1,2)$ for PCT and PCAT, and $(0.5,0.4)$ for RCAT, and retain these values for all the sequential updates. PCT significantly reduces the test error while containing NFs across the four updates, but clearly worsens the robustness and RNFs. PCAT, relying on adversarial training, mitigates this issue, but still exhibits NF and RNF rates around 3%. RCAT achieves the best trade-off again in this setting also, with comparable test and robust errors to those of PCAT, while reporting consistently lower NF and RNF rates of about 2%. This confirms the capability of RCAT to also deal with the case of multiple sequential updates.
TABLE IV: Multiple sequential updates on ImageNet models using PCT, PCAT, and RCAT. We show the evolution of test error, NFs, robust error, and RNFs (in %), when applying each method sequentially, i.e., using the newly-trained model as a baseline for the next update.
Test Error NFs Robust Error RNFs
PCT $(I_{1},I_{2})$ 35.50 2.28 74.30 6.14
$(I_{2},I_{3})$ 28.90 3.16 84.08 18.32
$(I_{3},I_{4})$ 22.88 2.68 69.72 15.30
$(I_{4},I_{5})$ 19.64 2.84 70.28 27.22
PCAT $(I_{1},I_{2})$ 38.62 2.58 65.12 2.20
$(I_{2},I_{3})$ 33.36 4.76 60.56 3.98
$(I_{3},I_{4})$ 25.78 2.96 44.42 1.94
$(I_{4},I_{5})$ 21.86 3.32 40.18 3.44
RCAT $(I_{1},I_{2})$ 37.98 1.84 64.76 1.96
$(I_{2},I_{3})$ 31.00 2.30 55.64 1.58
$(I_{3},I_{4})$ 23.86 2.30 43.28 1.66
$(I_{4},I_{5})$ 21.48 2.42 39.22 2.02
To summarize, performing model updates that retain high accuracy and robustness with low regression of both metrics is challenging. PCT tends to recover accuracy, at the expense of worsening robustness, as it was not originally designed to also deal with adversarial robustness. PCAT and RCAT find better accuracy-robustness trade-offs, with RCAT outperforming PCAT. In particular, RCAT provides comparable improvements in accuracy while preserving or even improving robustness and, at the same time, minimizing the sum of NFs and RNFs, thereby effectively reducing regression of both accuracy and robustness.